THE STUDY OF RARE GAS-COINAGE METAL HALIDE INTERACTIONS: FOURIER TRANSFORM MI-CROWAVE SPECTROSCOPY OF Kr-CuX (X=F,Cl)

\underline{JULIE M. MICHAUD}, STEPHEN A. COOKE AND MICHAEL C. L. GERRY, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, B.C., Canada V6T 1Z1.

The rotational spectra of KrCuX (X=F,Cl) have been observed using high resolution Fourier transform microwave spectroscopy. The molecules were made using laser-ablation of a copper rod in the presence of a halide precursor contained in krypton gas. These molecules were stabilized using a supersonic jet into the cavity of the spectrometer. The measurements allowed the determination of rotational constants and centrifugal distortion constants (B_{0}, D_{0}) to high precision. The different isotopomers that were observed permitted the determination of the Kr-Cu and Cu-X distances. The nuclear quadrupole coupling constants for Cu, Cl and ^{83}Kr were measured. The results have been supported by ab initio calculations.